TIME-RESOLVED IR-IR DOUBLE RESONANCE FOR THE ν_{4} and 2ν_{4} − ν_{4} VIBRATION-ROTATION TRANSITIONS OF CH_{3}F

YUSUKE OKABAYASHI, \underline{JIAN TANG}, YUKI MIYAMOTO, KENTAROU KAWAGUCHI, Graduate School of Natural Science and Technology, Okayama University, 3-1-1 Tsushima-Naka, Okayama 700-8530, Japan.

Time-resolved double resonance (TRDR) method has been used to study the rotational and vibrational relaxiation for a long time. However, the study of IR-IR TRDR in the 3μm region has been very limited due to the lack of powerful tunable IR lasers in this region. In the present study, we used a tunable pulsed OPO IR laser (6 mJ@6 ns, ∼0.2 cm^{−1} linewidth) and a tunable cw-OPO IR laser (up to 200 mW, ∼3 MHz linewidth) both in the 3μm region to study the IR-IR TRDR for the ν_{4} and 2ν_{4} − ν_{4} vibrational bands of CH_{3}F. The pulsed IR laser pumps CH_{3}F from the ground state to a certain rotational level (J, K) of v_{4} = 1, and the cw IR laser probes this (J, K) level by the 2ν_{4} − ν_{4} ladder transition. A time profile for the transition intensity of the probe line gives the T_{1} relaxiation time for the (J, K) level of v_{4} = 1. The pressure dependence of T_{1} for various (J, K) levels of v_{4} = 1 was obtained, which can be compared with the linewidth measurement^{a} for the ν_{4} band of CH_{3}F.

aA. G. Cartlidge and R. J. Butcher, J. Phys. B: At. Mol. Opt. Phys. \underline{23}, 2083 (1990).